# CCFC: Bridging Federated Clustering and Contrastive Learning

### v1

@@ Abstract
Federated clustering, an essential extension of centralized clustering for federated scenarios, enables multiple data-holding clients to collaboratively group data while keeping their data locally. In centralized scenarios, clustering driven by representation learning has made significant advancements in handling high-dimensional complex data. However, the combination of federated clustering and representation learning remains underexplored. To bridge this, we first tailor a cluster-contrastive model for learning clustering-friendly representations. Then, we harness this model as the foundation for proposing a new federated clustering method, named cluster-contrastive federated clustering (CCFC). Benefiting from representation learning, the clustering performance of CCFC even double those of the best baseline methods in some cases. Compared to the most related baseline, the benefit results in substantial NMI score improvements of up to 0.4155 on the most conspicuous case. Moreover, CCFC also shows superior performance in handling device failures from a practical viewpoint.
@@ 1 Introduction
Federated clustering, an essential extension of centralized clustering for federated scenarios, enables multiple data-holding clients to collaboratively group data while keeping their data locally [8, 27, 31, 30]. It often constitutes a significant initial step in many learning tasks, such as client-selection [10] and personalization [20, 6]. And it is natural to extend centralized clustering methodologies into federated scenarios. In centralized scenarios, clustering driven by representation learning has made significant advancements in handling high-dimensional complex data, particularly images [34]. However, the marriage of federated clustering and representation learning (e.g. contrastive learning [24, 2]) remains an underexplored research avenue.
@@
To illustrate the necessity of the marriage and the challenges it poses, we simulate a simple federated scenario on MNIST, where images are evenly distributed among 10 clients. Fig. 0(a) showcases the t-SNE visualization of a randomly selected local dataset in its original data space, and one can observe that many samples from different categories are mixed up together, necessitating representation learning. To handle this, there are two straightforward methods: 1) Independent execution of a SimSiam model [5] by each client; 2) Integration of the SimSiam model into the Federated Averaging (FedAvg) framework [21]. As shown in Fig. 0(b) and Fig. 0(c), one can see that: 1) The local model trained solely on local data exhibits poor performance. 2) The collaborative training yields a superior global model. However, in the subsequent round of local updates, the limited information from the local data can lead to model regression (Fig. 0(d)). A similar problem of the model regression has also been observed in federated classification tasks, and it can be alleviated by introducing a model-contrastive loss to regularize the local model to refrain from deviating excessively from the global model during the local training [17].
Motivated by these, we first tailor a contrastive model, named cluster-contrastive model, for learning clustering-friendly representations. Then, we harness this model as the foundation for proposing a simple but effective federated clustering method, named cluster-contrastive federated clustering (CCFC), which embeds the cluster-contrastive model into the FedAvg framework with the model-contrastive regularization term. Throughout the entire training procedure of CCFC, the only shared information between clients and the server is cluster-contrastive models and cluster centroids, thereby preserving data privacy. In each communication round, CCFC involves three main steps: global information dissemination, local training and local information aggregation. In the first step, each client first downloads a global model and $k$ global cluster centroids from the server, and then update the local model with the downloaded one, where $k$ is the number of clusters. In the second step, each client performs cluster assignment by labeling their local data with the index of the nearest global cluster centroid in the latent space, and they train the local model using both the local data and the assignment results. After this, k-means (KM) [19] is employed to create $k$ local cluster centroids for storing local semantic information. In the final step, clients upload their trained local models and local cluster centroids to the server, which aggregates these local models into a new global model through weighted averaging, and aggregates these local cluster centroids into $k$ new global cluster centroids through KM.
Comprehensive experiments demonstrate the superiority of the cluster-contrastive model in learning clustering-friendly representations, as well as the excellence of the proposed clustering method in terms of clustering performance and handling device failures. The most related method to our CCFC is k-fed [8], with the main distinction being the inclusion of a representation learning process in CCFC. Surprisingly, this learning process results in substantial NMI [28] score improvements of up to 0.4155 and Kappa [18] score enhancements of 0.4593 on the most conspicuous case. In summary, our contributions are as follows:
2 Related Work
2.1 Clustering
To address these, federated clustering (FC) has arisen, allowing multiple clients to collaboratively group data while preserving their data locally [8, 27, 31, 30]. As an extension of centralized clustering, FC inherently embodies an exploratory path, i.e. extending centralized clustering methodologies to federated scenarios. Representative extensions include k-FED [8] and SDA-FC-KM [31], derived from k-means [19]; FFCM [27] and SDA-FC-FCM [31], extensions of fuzzy c-means [3]; and PPFC-GAN [30], an outgrowth of DCN [32]. Despite these extensions have advanced FC, their progress is limited. For instance, on the MNIST dataset, the NMI [28] values of the top 10 centralized clustering methods surpass 0.93111https://paperswithcode.com/sota/image-clustering-on-mnist-full. In stark contrast, even in the simplest federated scenarios, where local data among clients is independently-and-identically-distributed (IID), the NMI value of the cutting-edge PPFC-GAN falls short of 0.7 (Tab. 1).
2.2 Contrastive Learning
Contrastive learning [24, 2] has emerged as a prominent paradigm within the realm of representation learning, particularly for the acquisition of visual representations. While contrastive learning has been extensively explored and proven effective in centralized scenarios [9, 26, 22], its extension into federated scenarios has been relatively limited. These extensions can be broadly categorized into two groups: one employs contrastive learning to handle the challenges posed by the non-IID problem in federated classification tasks [17, 29, 23], while the other harnesses it for the acquisition of generic representations for downstream learning tasks [35, 12, 36].
As far as we can ascertain, no extensions have been introduced to tackle federated clustering tasks. To bridge contrastive learning and federated clustering, we tailor a new contrastive model for clustering tasks, called cluster-contrastive model. Leveraging this model, we propose a simple yet effective federated deep clustering method named cluster-contrastive federated clustering (CCFC).
3 Cluster-Contrastive Federated Clustering
3.1 Cluster-Contrastive Model
Given $k$ clusters obtained from clustering, each one comprises $n_{c}$ images. As shown in Fig. 2, the proposed model takes $n_{c}$ images from cluster $c$ as inputs. These images are processed by the same encoder $f$, which comprises a backbone (e.g., ResNet-18 [13]) and an MLP projector [4]. The latent representations of these images are denoted as $z_{i,\,c}=f(x_{i,\,c}),\,i=1,\,\cdots,\,n_{c}$. Then, an MLP predictor [11], denoted as $h$, transforms the $i$-th latent representation and matches it to the remaining latent representations. The prediction of $x_{i}$ for the remaining images is denoted as $p_{i,\,c}=h(f(x_{i,\,c}))$. The negative average cosine similarity of $x_{i,\,c}$ to other images is defined as:
$\mathcal{L}=\frac{1}{kn_{c}}\sum_{c=1}^{k}\sum_{i=1}^{n_{c}}D\left(x_{i,\,c},\,stopgrad\left(\{x_{j,\,c}\}_{j\neq i}^{n_{c}}\right)\right),$ (2)
Actually, the proposed model is a more general form of simple Siamese (SimSiam) networks [5]. Consider a scenario where $k=n$ ($n$ is the number of images) and $n_{c}=2$, signifying that each image forms a cluster, with each cluster consists of the two augmented views of that image, the proposed model reverts to SimSiam. In this context, the proposed model leans towards grasping low-level semantics through the utilization of sample-level disparities among images, and the resulting representations are suboptimal for clustering tasks. In contrast, the latent representation learned by the generalized SimSiam (cluster-contrastive model), which takes into account the cluster-level distinctions among images, is more clustering-friendly.
3.2 Cluster-Contrastive Federated Clustering
Given a real-world dataset $X$ distributed among $m$ clients, i.e., $X=\bigcup_{l=1}^{m}X^{l}$. There are two straightforward methods for conducting contrastive learning in federated scenarios here: 1) Independent execution of contrastive learning by each client; 2) Integration of contrastive learning into the Federated Averaging (FedAvg) [21] framework. However, their performance is unsatisfactory. As shown in Fig. 1, one can observe: 1) The local model trained solely on local data exhibits poor performance. 2) Although the collaborative training can yield a better-performing global model, in the subsequent round of local updates, the limited information from the local data can lead to model regression. A similar problem of the model regression has also been observed in federated classification tasks, and it can be alleviated by introducing a model-contrastive loss to regularize the local model to refrain from deviating excessively from the global model during the local training [17].
@@ 3.2 Cluster-Contrastive Federated Clustering
Motivated by these, we propose a new federated clustering model, named as cluster-contrastive federated clustering (CCFC), which embeds the cluster-contrastive model into the FedAvg framework with the model-contrastive regularization term. As shown in Fig. 3 and Algorithm 1, CCFC involves three steps in each communication round: global information dissemination, local training and local information aggregation. Details are given below.
@@
Then, each client $l$ trains the local model $w^{l}$ using their local data $X^{l}$ and the assignment results. To alleviate the model regression problem, we introduce a model-contrastive regularization term into Eq. 2. Finally, the revised loss function is denoted as:
$\mathcal{L_{R}}=\mathcal{L}+\frac{\lambda}{kn_{c}}\sum_{c=1}^{k}\sum_{i=1}^{n_{c}}R(w^{l}(x_{i,\,c}),\,w^{g}(x_{i,\,c})),$ (4)
@@
and $\lambda$ is the tradeoff hyperparameter. By minimizing Eq. 4, the first term will encourage the local model to encode the local semantic structures discovered by the assignment result into the latent representation space, while the second one will encourage the local model $w^{l}$ to refrain from deviating excessively from the global model $w^{g}$. During the local training, the parameters of the global model remain fixed, while only those of the local model undergo updates.
@@
So far, the whole communication loop of CCFC has been built up. And the final clustering result can be obtained by solving Eq. 3.
@@ 8             Local training:
9                 group $X^{l}$ by solving Eq. 3
10                 train $w^{l}$ by minimizing Eq. 4
@@ 18            Local information aggregation:
Cluster assignment: solve Eq. 3

### v2

@@ Abstract
Federated clustering, an essential extension of centralized clustering for federated scenarios, enables multiple data-holding clients to collaboratively group data while keeping their data locally. In centralized scenarios, clustering driven by representation learning has made significant advancements in handling high-dimensional complex data. However, the combination of federated clustering and representation learning remains underexplored. To bridge this, we first tailor a cluster-contrastive model for learning clustering-friendly representations. Then, we harness this model as the foundation for proposing a new federated clustering method, named cluster-contrastive federated clustering (CCFC). Benefiting from representation learning, the clustering performance of CCFC even double those of the best baseline methods in some cases. Compared to the most related baseline, the benefit results in substantial NMI score improvements of up to 0.4155 on the most conspicuous case. Moreover, CCFC also shows superior performance in handling device failures from a practical viewpoint. Code is available at https://github.com/Jarvisyan/CCFC-pytorch.
keywords:
Federated clustering, representation learning, contrastive learning.
@@ 1 Introduction
Federated clustering (FC), an essential extension of centralized clustering for federated scenarios, enables multiple data-holding clients to collaboratively group data while keeping their data locally [1, 2, 3, 4]. It often constitutes a significant initial step in many learning tasks, such as client-selection [5] and personalization [6, 7].
As an extension of centralized clustering, it is natural to extend centralized clustering methodologies into FC. For instance, k-fed [1] is an extension of k-means (KM) [8], while federated fuzzy c-means (FFCM) [2] builds upon fuzzy c-means (FCM) [9]. Both of these extended approaches employ a similar methodology: alternately estimating local and global cluster centroids, and labeling the local data based on their proximity to these global centroids. Nevertheless, the global cluster centroids are vulnerable to heterogeneous client data, diminishing both the robustness and overall performance of the model [3]. To tackle this, synthetic data aided federated clustering (SDA-FC) [3] and privacy-preserving federated deep clustering based on GAN (PPFC-GAN) [4] bridged FC and deep generative model, enabling the application of centralized clustering techniques to synthetic data, and safeguarding global cluster centroids from data heterogeneity problem. Although these methods have advanced FC, a substantial gap persists between FC and centralized clustering. For instance, on the MNIST dataset, the NMI [10] values of the top 10 centralized clustering methods surpass 0.93111https://paperswithcode.com/sota/image-clustering-on-mnist-full. In stark contrast, even in the simplest federated scenario, where local data among clients is independently-and-identically-distributed (IID), the NMI value of the cutting-edge PPFC-GAN falls short of 0.7 (Table 1).
To bridge this gap, it is logical to revisit centralized clustering methodologies, especially the most advanced ones. In centralized scenarios, clustering driven by representation learning has made significant advancements in handling high-dimensional complex data, particularly images [11, 12]. However, the marriage of federated clustering and representation learning (e.g. contrastive learning [13, 14]) remains an underexplored research avenue. As PPFC-GAN has validated the feasibility of training deep learning model using shared synthetic data, a obvious next step is to explore contrastive learning on this data. However, the increased utility of synthetic data often comes with a heightened risk of deep data leakage from shared synthetic data [15]. In contrast to synthetic data, which may serve as replicas of real data and thus pose a direct privacy risk [16], sharing model parameters offers a significantly greater degree of privacy preservation. Recall that the core distinction between FC and centralized clustering lies in privacy preservation. Hence, we choose to use the classical FedAvg framework [17] to bridge FC and contrastive learning, rather than synthetic data aided framework [3, 4].
Motivated by these, we first tailor a contrastive model, named cluster-contrastive model, for learning clustering-friendly representations that show improved intra-cluster compactness and inter-cluster separability. Then, we harness this model as the foundation for proposing a simple but effective federated clustering method, named cluster-contrastive federated clustering (CCFC), which embeds the cluster-contrastive model into the FedAvg framework. Throughout the entire training procedure of CCFC, the only shared information between clients and the server is cluster-contrastive models and cluster centroids, thereby preserving data privacy. As shown in Figure 1, CCFC involves three main steps in each communication round: global information dissemination, local training and local information aggregation. In the first step, each client first downloads a global model and $k$ global cluster centroids from the server, and then update the local model with the downloaded one, where $k$ is the number of clusters. In the second step, each client performs cluster assignment by labeling their local data with the index of the nearest global cluster centroid in the latent space, and they train the local model using both the local data and the assignment results. After this, KM is employed to create $k$ local cluster centroids for storing local semantic information. In the final step, clients upload their trained local models and local cluster centroids to the server, which aggregates these local models into a new global model through weighted averaging, and aggregates these local cluster centroids into $k$ new global cluster centroids through KM.
As widely acknowledged, the performance of the FedAvg framework can be hindered by the issue of data heterogeneity, and this problem remains unresolved and challenging [15, 18, 19]. Thus, this work focuses on how contrastive learning aids in FC and attempts to analyze how data heterogeneity impairs the clustering-friendliness of the learned representations, while addressing this issue of data heterogeneity is deferred to future work. Comprehensive experiments demonstrate the superiority of the cluster-contrastive model in learning clustering-friendly representations, as well as the excellence of the proposed clustering method in terms of clustering performance and handling device failures. The most related method to our CCFC is k-fed [1], with the main distinction being the inclusion of a representation learning process in CCFC. Surprisingly, this learning process results in substantial NMI [10] score improvements of up to 0.4155 and Kappa [20] score enhancements of 0.4593 on the most conspicuous case. With the rise of data heterogeneity, the learned representations progressively lose their clustering-friendly traits, as evidenced by semantic inconsistency in the latent spaces of local models and overlapping cluster representations (cluster collapse).
In summary, our contributions are as follows:
  • 1.
  • 2.
  • 3.
  • 4.
    A systematic analysis of how data heterogeneity impairs the clustering-friendliness of the learned representations offers a potential direction for alleviating the data heterogeneity problem.
The remainder of this paper is structured as follows: Section 2 provides a concise review of representative federated clustering methods and discusses the current state of contrastive learning research in federated scenarios. Section 3 presents a new cluster-contrastive model, upon which the CCFC, a simple yet effective federated clustering method, is developed. Section 4 highlights the advantages of the proposed methods and conducts a comprehensive analysis of how data heterogeneity impairs the clustering-friendliness of the learned representations. Finally, Section 5 concludes the paper.
2 Related Work
In this section, we first provide a concise review of representative federated clustering methods and then discuss the current state of contrastive learning research in federated scenarios.
2.1 Clustering
To address these, federated clustering (FC) has arisen, allowing multiple clients to collaboratively group data while preserving their data locally [1, 2, 3, 4]. As an extension of centralized clustering, FC inherently embodies an exploratory path, i.e. extending centralized clustering methodologies to federated scenarios. Representative extensions include k-FED [1] and SDA-FC-KM [3], derived from k-means (KM) [8]; FFCM [2] and SDA-FC-FCM [3], extensions of fuzzy c-means (FCM) [9]; and PPFC-GAN [4], an outgrowth of DCN [25]. The key idea behind these extensions is similar: constructing $k$ global cluster centroids and labeling local data with the index of the nearest global centroid, where $k$ is the number of clusters. In methods k-FED and FFCM, each client first employs KM or FCM to extract local cluster centroids from their local data and then transmit these centroids to the server, where these local centroids are merged into $k$ global centroids through KM. Nonetheless, the global cluster centroids are fragile and sensitive to heterogeneous data among clients, which undermines the model’s robustness and model performance. To address this, SDA-FC-KM, SDA-FC-FCM, and PPFC-GAN bridged FC and deep generative model, enabling the application of centralized clustering techniques to synthetic data for the construction of $k$ global cluster centroids. Although these extensions have advanced FC, a substantial gap persists between FC and centralized clustering. For instance, on the MNIST dataset, the NMI values of the top 10 centralized clustering methods surpass 0.93222https://paperswithcode.com/sota/image-clustering-on-mnist-full. In stark contrast, even in the simplest federated scenario, where local data among clients is independently-and-identically-distributed (IID), the NMI value of the cutting-edge PPFC-GAN falls short of 0.7 (Table 1).
2.2 Contrastive Learning
Contrastive learning [13, 14] has emerged as a prominent paradigm within the realm of representation learning, particularly for the acquisition of visual representations. In early studies [26, 27], these methods aims to maximize the similarity between augmented views of the same image (i.e., positive pairs) while minimizing the similarity between augmented views of different images (i.e., negative pairs). In these methods, the contrast between the negative pairs is considered a key factor in preventing model collapse (i.e., the model possesses a trivial solution where all outputs collapse to a constant). Nevertheless, further studies [28, 29] have shown that the negative pairs are not essential for averting model collapse. While contrastive learning has been extensively explored and proven effective in centralized scenarios [30, 31, 32], its extension into federated scenarios has been relatively limited. These extensions can be broadly categorized into two groups: one employs contrastive learning to handle the challenges posed by the non-IID problem in federated classification tasks [33, 34, 35], while the other harnesses it for the acquisition of generic representations for downstream learning tasks [36, 37, 38].
In FC, there is a concurrent work called federated momentum contrastive clustering (FedMCC) [39], which is an extension of contrastive clustering (CC) [40]. CC is a centralized clustering approach that explicitly executes contrastive learning at both instance and cluster levels by maximizing positive pair similarities and minimizing negative pair similarities in the row and column representation spaces, respectively. Note that the objective function of CC, besides contrastive loss terms, incorporates an entropy-based regularization term to prevent trivial solutions and encourage uniform cluster sizes [40, 41]. This technique to avoid trivial solutions has been carried over in FedMCC. Although many real-world datasets in centralized scenarios, like Fashion-MNIST, CIFAR-10 [42], and STL-10 [43], feature a uniform class distribution, the training data distribution in federated environments is generally imbalanced and heterogeneous due to variations in client preferences [44, 45]. Consequently, encouraging uniform cluster sizes in FedMCC becomes unreasonable, thus restricting its real-world applicability.
By contrast, our study introduces a new contrastive model for clustering tasks, called cluster-contrastive model, which places no limitations on cluster sizes, thus offering greater flexibility. Leveraging this model, we further propose a simple yet effective federated deep clustering method named cluster-contrastive federated clustering (CCFC).
3 Cluster-Contrastive Federated Clustering
In this section, we first propose a cluster-contrastive model to encode the semantic structures discovered by clustering into the latent representation space. Then, we harness this model as the foundation for proposing a simple but effective federated clustering method.
3.1 Cluster-Contrastive Model
In this work, due to the elegance of the SimSiam [29], we select it as the prototype model among numerous contrastive models. It maximizes the similarity between augmented views of the same image (i.e., positive pairs) directly, without minimizing the similarity between augmented views of different images (i.e., negative pairs) or using a momentum encoder. Contrasting negative pairs usually requires some customized techniques (such as memory bank [46] and memory queue [27]) to perform large-batch training. The momentum encoder [28] introduces extra hyperparameter and necessitates additional memory allocation for storing historical momentum information. Both of these can complicate the model training and incur extra computational costs. However, the representations learned by SimSiam may be suboptimal for clustering tasks, as it leans towards grasping low-level semantics through sample-level differences among images, neglecting the cluster-level distinctions among them.
To handle this, we tailor a new contrastive model for clustering tasks by extending SimSiam, called cluster-contrastive model. It maximizes the similarity among images within the same cluster to learn cluster-invariant representations, i.e. samples within the same cluster are expected to have similar latent representations. Formally, given $k$ clusters obtained from clustering, each one comprises $n_{c}$ images. As shown in Figure 2, the proposed model takes $n_{c}$ images from cluster $c$ as inputs. These images are processed by the same encoder $f$, which comprises a backbone (e.g., ResNet-18 [47]) and an MLP projector [26]. The latent representations of these images are denoted as $z_{i,\,c}=f(x_{i,\,c}),\,i=1,\,\cdots,\,n_{c}$. Then, an MLP predictor [28], denoted as $h$, transforms the $i$-th latent representation and matches it to the remaining latent representations. The prediction of $x_{i}$ for the remaining images is denoted as $p_{i,\,c}=h(f(x_{i,\,c}))$. The negative average cosine similarity of $x_{i,\,c}$ to other images is defined as:
$\mathcal{L}=\sum_{c=1}^{k}\frac{1}{kn_{c}}\sum_{i=1}^{n_{c}}D\left(x_{i,\,c},\,stopgrad\left(\{x_{j,\,c}\}_{j\neq i}^{n_{c}}\right)\right),$ (2)
3.2 Cluster-Contrastive Federated Clustering
Given a real-world dataset $X$ distributed among $m$ clients, i.e., $X=\bigcup_{l=1}^{m}X^{l}$. There are two straightforward methods for integrating FC with contrastive learning here: 1) Independent execution of contrastive learning by each client; 2) Integration of contrastive learning into the Federated Averaging (FedAvg) [17] framework. To illustrate the necessity of the integration and the challenges it poses, we simulate a simple federated scenario on MNIST, where images are evenly distributed among 10 clients. Figure 3(a) showcases the t-SNE visualization of a randomly selected local dataset in its original data space, and one can observe that many samples from different categories are mixed up together, necessitating representation learning. As shown in Figure 3(b) and Figure 3(c), one can see that: 1) The local model trained solely on local data exhibits poor performance. 2) The collaborative training yields a superior global model. However, in the subsequent round of local updates, the limited information from the local data can lead to model regression (Figure 3(d)). A similar problem of the model regression has also been observed in federated classification tasks, and it can be alleviated by introducing a model-contrastive loss to regularize the local model to refrain from deviating excessively from the global model during the local training [33].
@@ 3.2 Cluster-Contrastive Federated Clustering
Motivated by these, we propose a new federated clustering model, named as cluster-contrastive federated clustering (CCFC), which embeds the cluster-contrastive model into the FedAvg framework with the model-contrastive regularization term. As shown in Figure 1 and Algorithm 1, CCFC involves three steps in each communication round: global information dissemination, local training and local information aggregation. Details are given below.
@@
Then, each client $l$ trains the local model $w^{l}$ using their local data $X^{l}$ and the assignment results. To alleviate the model regression problem, we introduce a model-contrastive regularization term into Equation 2. Finally, the revised loss function is denoted as:
$\mathcal{L_{R}}=\mathcal{L}+\sum_{c=1}^{k}\frac{\lambda}{kn_{c}}\sum_{i=1}^{n_{c}}R(w^{l}(x_{i,\,c}),\,w^{g}(x_{i,\,c})),$ (4)
@@
and $\lambda$ is the tradeoff hyperparameter. By minimizing Equation 4, the first term will encourage the local model to encode the local semantic structures discovered by the assignment result into the latent representation space. Note that the parameters of the global model $w^{g}$ remain fixed throughout the local training, whereas only those of the local model $w^{l}$ undergo updates. Consequently, the second term will correct the local updates by maximizing the agreement (cosine similarity) of representation learned by the current local model $w^{l}$ and that learned by the global model $w^{g}$ from the previous communication round.
@@
So far, the whole communication loop of CCFC has been built up. And the final clustering result can be obtained by solving Equation 3.
@@ 8             Local training:
9                 group $X^{l}$ by solving Equation 3
10                 train $w^{l}$ by minimizing Equation 4
@@ 18            Local information aggregation:
Cluster assignment: solve Equation 3
@@
4 Experiment
In this section, we detail the experimental settings, assesses the effectiveness of our proposed methods, analyze the clustering-friendliness of the learned representations in different federated scenarios, and conduct a sensitivity analysis of clustering methods to device failures. Finally, we present a summary of the experimental findings.
4.1 Experimental Setup
The proposed method is trained using the Adam optimizer [49], implemented in the PyTorch [50].